OpenStudy (anonymous):

(c) Find the equations of the curves whose parametric equations are: ii) x = 3cos theta , y = 3sin theta

OpenStudy (anonymous):

The aim is to eliminate the parameter theta....the final eqn should be a relationship between x and y

OpenStudy (anonymous):

x = 3cos t , y = 3sin t

OpenStudy (anonymous):

what relationship between sin t and cos t exists

OpenStudy (anonymous):

to sole t is write instead of theta ?

hartnn (hartnn):

if you square and add both the equations, you will eliminate theta

OpenStudy (anonymous):

x^2 = 9 cos t^2 , y = 9 sin t^ 2 ???

hartnn (hartnn):

its y^2 = 9 sin t^2 and yes, now add them, what u get ?

OpenStudy (anonymous):

x^2 = 9 cos t^2 , y^2 = 9 sin t^2

OpenStudy (anonymous):

solve them together or spread

OpenStudy (anonymous):

i don"t speak eng good :(

hartnn (hartnn):

don't forget to put brackets x^2 = 9 (cos t)^2 , y^2 = 9 (sin t)^2 when you add them, u get x^2+y^2 = 9[sin^2 t + cos^2 t] u know what [sin^2 t + cos^2 t] equals ?

OpenStudy (anonymous):

i don"t sure

hartnn (hartnn):

[sin^2 t + cos^2 t] equals 1 so, finally u get x^2+y^2 = 9

OpenStudy (anonymous):

Thank you you're a genius Hahahah

hartnn (hartnn):

\[ \begin{array}l\color{red}{\text{w}}\color{orange}{\text{e}}\color{#e6e600}{\text{l}}\color{green}{\text{c}}\color{blue}{\text{o}}\color{purple}{\text{m}}\color{purple}{\text{e}}\color{red}{\text{ }}\color{orange}{\text{^}}\color{#e6e600}{\text{_}}\color{green}{\text{^}}\color{blue}{\text{}}\end{array} \]

OpenStudy (anonymous):

Thank you :)